### v2

@@ 4.2 Comparison with SOTA
Table 1: Quantitative Comparison on four large-scale scene datasets. The ’-’ symbol indicates Mega-NeRF [33] and Switch-NeRF [46] were not evaluated on MatrixCity due to difficulties in adjusting its training configurations beyond the provided, resulting in poor performance on this dataset. The best results of each metric are in bold.
@@
MegaNeRF [33] - - - 0.628 22.08 0.489 0.553 24.06 0.516 0.547 20.93 0.504
Switch-NeRF [46] - - - 0.654 22.57 0.457 0.562 24.31 0.496 0.579 21.54 0.474
GP-NeRF [40] 0.611 23.56 0.630 0.661 22.31 0.448 0.565 24.06 0.496 0.566 21.03 0.486
${}^{\dagger}$ [9] 0.735 23.67 0.384 0.791 21.44 0.236 0.777 25.47 0.2774 0.720 20.46 0.305
Ours 0.865 27.46 0.204 0.813 22.00 0.211 0.813 25.77 0.228 0.778 21.55 0.246
For an apple-to-apple rendering quality comparison with the SOTA large-scale scene reconstruction strategy, we use the performance of the CityGS with no LoD. The quantitative result is shown in Table 1. Due to page limits, we put results of Sci-Art in the Appendix. It can be observed that our method outperforms the NeRF-based baselines by a large margin. On the one hand, as far as we know, our method is the first of all attempts that successfully reconstructs the whole MatrixCity with highly variable camera altitude, ranging from 150m to 500m [13]. The PSNR arrives at 27.46 and qualitative results presented in Fig. 5 also validate the high fidelity of our renderings. Compared with the baseline ${}^{\dagger}$, our CityGS can capture much richer details. More visualizations on MatrixCity can be found in the Appendix. On the other three realistic scenes, our method enables much higher SSIM and LPIPS, which indicates outstanding visual quality. As shown in Fig. 4, thin structures like girder steel and window frames can be well reconstructed. The details of complex structures like grass and rubble are well recovered as well.
@@ 4.4 Ablation
Table 3: Ablation on block numbers and SSIM threshold $\varepsilon$. The experiment is conducted on the Rubble dataset. The first row is the performance of coarse global Gaussians prior mentioned in Sec. 3.2.1, and thus has no block number or $\varepsilon$ setting. MEAN denotes an average number of assigned training poses among divided blocks. The best performances are in bold
@@
n/a n/a 0.770 24.83 0.284 n/a
0.1 $2\times 2$ 0.798 25.53 0.250 835
0.1 $3\times 3$ 0.809 25.75 0.239 552
0.1 $4\times 4$ 0.809 25.62 0.233 432
0.12 $3\times 3$ 0.813 25.77 0.231 469
0.14 $3\times 3$ 0.812 25.43 0.229 415
In this section, we first explore the influence of hyper-parameters in training, namely block number and data assignment threshold $\varepsilon$ mentioned in Sec. 3.2. Comparing the first row in Table 3 with others, it can be observed that the proposed divide-and-conquer training strategy effectively improves from coarse global Gaussian prior in all metrics. On the other hand, as shown in Table 3, as the block number grows, the average data assigned to blocks decreases. As the total iteration of block-wise finetuning is fixed, the regional Gaussians gain more chances to refine details, leading to better SSIM and LPIPS. The $\varepsilon$ also controls data assignment. As $\varepsilon$ grows, the average poses assigned decreases. And if it is too high, many necessary training data will be lost, thus leading to lower PSNR performance.
@@
The quantitative ablation on LoD strategy is shown in Table 4. Comparing the first and the second row, the computation burden introduced by the point-wise strategy leads to the considerably worse real-time performance. Its rendering quality is on par with the smaller interval setting, i.e. third row. Comparing the first row with the second and the third row, it can be observed that with larger intervals for LoD 2 and LoD 1, the rendered results would contain richer details, but downgraded speed. And first row arrives at a balance, which is adopted as the standard setting.
4.5 Scene Manipulation
Figure 7: Illustration of city scene manipulation driven by explicit representation of CityGS. In Part (a), the dome of the original building in the first row is repainted to the desired color shown in the second row. In Part (b), the building of the first row is removed and replaced with the one shown in the second row. In Part (c), the cars parked at locations shown in the left image are moved to the positions shown in the right image, so as to simulate the required traffic conditions. NeRF-based methods struggle to realize such manipulation.
For the implicit representation of NeRF-based methods, it is hard to explain the correspondence between network parameters and scene structure. However, since we can reconstruct the explicit city representation with relatively high geometric precision in CityGS, the geometric and appearance distribution can be manipulated as desired. The demos are shown in Fig. 7. The appearance of a specified part of a building can be transformed to the desired style. It is also possible to delete a building and replace it with another one. By placing cars or pedestrians, the pre-defined traffic conditions can be simulated in the city. These demos indicate potential real-time and interactive application of CityGS.
@@ A Additional Experimental Results
Besides, as shown in Table S1, LoD significantly improves efficiency, especially for extremely large-scale scenes such as MatrixCity. Compared with our CityGS, the ${}^{\dagger}$ [9] possesses faster speed but significantly lower rendering quality. The main reason is that the original 3DGS requires sufficiently large iterations and memory to optimize the whole scene with thousands of images. Bounded by computation resources, the capacity of the trained original 3DGS is too limited to well represent the whole large-scale scene.
@@ B Detailed Parameter Setting
We present the specific hyper-parameter configurations for each dataset in Table S2. The roles of training parameters are detailed in Sec. 3.2, while the role of rendering parameter Distance Interval is specified in Sec. 3.3. Note that for LoD on datasets except for MatrixCity, we use three detail levels of compression rate 60%, 50%, and 40%.
@@ D Qualitative Validation on Concatenated Fusion
To validate the effectiveness of the concatenated fusion strategy, we perform rendering at the viewpoints where the visible area spans multiple blocks. The Gaussians utilized here come from direct concatenation of fine-tuned Gaussians of corresponding blocks. As depicted in Fig. S2, rendering from a specific viewpoint may involve four or more blocks. Despite that, the rendered images exhibit no discernible discontinuities, showcasing smooth boundary transitions facilitated by our coarse global Gaussian prior, as discussed in Sec. 3.2.

### v3

@@ 4.2 Comparison with SOTA
For an apple-to-apple rendering quality comparison with the SOTA large-scale scene reconstruction strategy, we use the performance of the CityGS with no LoD. The quantitative result is shown in Table 1. Due to page limits, we put results of Sci-Art in the Appendix. It can be observed that our method outperforms the NeRF-based baselines by a large margin. On the one hand, as far as we know, our method is the first of all attempts that successfully reconstructs the whole MatrixCity with highly variable camera altitude, ranging from 150m to 500m [16]. The PSNR arrives at 27.46 and qualitative results presented in Fig. 5 also validate the high fidelity of our renderings. Compared with the baseline 3DGS, our CityGS can capture much richer details. More visualizations on MatrixCity can be found in the Appendix. On the other three realistic scenes, our method enables much higher SSIM and LPIPS, which indicates outstanding visual quality. As shown in Fig. 4, thin structures like girder steel and window frames can be well reconstructed. The details of complex structures like grass and rubble are well recovered as well. In addition, results in Table 3 validates that our training strategy can generalize well to street scenes.
Table 1: Quantitative Comparison on four large-scale scene datasets. The ’-’ symbol indicates Mega-NeRF [36] and Switch-NeRF [49] were not evaluated on MatrixCity due to difficulties in adjusting its training configurations beyond the provided, resulting in poor performance on this dataset. ’n/a’ means the number of Guassians ’#GS’ does not apply to NeRF-based methods. The best results of each metric are in bold.
@@
MegaNeRF [36] - - - n/a 0.628 22.08 0.489 n/a 0.553 24.06 0.516 n/a 0.547 20.93 0.504 n/a
Switch-NeRF [49] - - - n/a 0.654 22.57 0.457 n/a 0.562 24.31 0.496 n/a 0.579 21.54 0.474 n/a
GP-NeRF [43] 0.611 23.56 0.630 n/a 0.661 22.31 0.448 n/a 0.565 24.06 0.496 n/a 0.566 21.03 0.486 n/a
3DGS [12] 0.735 23.67 0.384 9.7M 0.791 21.44 0.236 6.6M 0.777 25.47 0.2774 6.1M 0.720 20.46 0.305 6.4M
Ours 0.865 27.46 0.204 23.7M 0.813 22.00 0.211 10.8M 0.813 25.77 0.228 9.7M 0.778 21.55 0.246 13.2M
@@ 4.3 Level of Detail
Table 3: Quantitative Comparison on MatrixCity street scene Block_A. The best results of each metric are in bold.
[…]
MipNeRF360[3] 0.717 22.00 0.488
3DGS[12] 0.701 21.14 0.441
CityGS 0.808 22.98 0.301
@@ 4.4 Ablation
Table 4: Ablation on training strategy. The experiment is conducted on the Rubble dataset. The baseline denotes simply selecting cameras within the 1.5× larger area of the block for training. The best performances are in bold.
@@
baseline 0.779 23.98 0.251 12.2
w global, baseline 0.801 25.01 0.227 15.4
w/o Eq.(3), Ours 0.749 23.43 0.252 19.7
w/o Eq.(4), Ours 0.810 25.45 0.233 9.6
Full, Ours 0.813 25.77 0.231 9.7
In this section, we analyze the impact of training strategies as shown in Table 4. As a baseline, we trained using cameras in a 1.5× larger area of each block. The results in the first and last lines of Table 4 indicate that this method significantly underperforms compared to our CityGS. The main issue is floaters caused by overfitting areas outside the block, but this is mitigated with guidance from the global model, improving performance as seen in the second line of Table 4. However, this improved baseline uses 1.5× more Gaussians than CityGS, making it inefficient. The last three lines demonstrate that Eq.(3) is crucial for data partitioning, and we found Eq.(4) helps prevent floaters at block edges. Further hyper-parameter ablations are detailed in the Appendix.
@@
The quantitative ablation on LoD strategy is shown in Table 5. Comparing the first and the second row, the computation burden introduced by the point-wise strategy leads to the considerably worse real-time performance. Its rendering quality is on par with the smaller interval setting, i.e. third row. Comparing the first row with the third and the fourth row, it can be observed that with larger intervals for LoD 2 and LoD 1, the rendered results would contain richer details, but downgraded speed. And first row arrives at a balance, which is adopted as the standard setting.
@@ 5 Conclusions
Acknowledgements
This work was supported in part by the National Key R&D Program of China (No. 2022ZD0116500), the National Natural Science Foundation of China (No. U21B2042, No. 62320106010), and in part by the 2035 Innovation Program of CAS, and the InnoHK program.
@@ A.1 Additional Quantitative Comparison
Table S2: Ablation on block numbers and SSIM threshold $\varepsilon$. The experiment is conducted on the Rubble dataset. The first row is the performance of coarse global Gaussians prior mentioned in Sec. 3.2 of main paper, and thus has no block number or $\varepsilon$ setting. MEAN denotes an average number of assigned training poses among divided blocks. The best performances are in bold.
[…]
0.1 $2\times 2$ 0.762 24.48 0.285 835
0.1 $3\times 3$ 0.806 25.41 0.238 552
0.1 $4\times 4$ 0.804 25.45 0.241 432
0.12 $3\times 3$ 0.813 25.77 0.231 469
0.14 $3\times 3$ 0.812 25.43 0.229 415
Besides, as shown in Table S1, LoD significantly improves efficiency, especially for extremely large-scale scenes such as MatrixCity. Compared with our CityGS, the 3DGS [12] possesses faster speed but significantly lower rendering quality. The main reason is that the original 3DGS requires sufficiently large iterations and memory to optimize the whole scene with thousands of images. Bounded by computation resources, the capacity of the trained original 3DGS is too limited to represent the whole large-scale scene well.
A.2 Additional Ablations
We also explored the influence of hyper-parameters in training, namely block number and data assignment threshold $\varepsilon$ mentioned in Sec. 3.2 of main paper. Here, we control the overall finetuning iterations to be $9\times 32000$. As shown in Table S2, as the block number grows, the average data assigned to blocks decreases. We achieve optimal results around 3 × 3 partitions. The $\varepsilon$ also controls data assignment. As $\varepsilon$ grows, the average poses assigned decreases. And if it is too high, many necessary training data will be lost, thus leading to lower PSNR performance.
@@ B Detailed Parameter Setting
We present the specific hyper-parameter configurations for each dataset in Table S3. The roles of training parameters are detailed in Sec. 3.2 of the main paper, while the role of rendering parameter Distance Interval is specified in Sec. 3.3 of the main paper. Note that for LoD on datasets except for MatrixCity, we use three detail levels of compression rate 60%, 50%, and 40%.
@@ D Qualitative Validation on Concatenated Fusion
To validate the effectiveness of the concatenated fusion strategy, we perform rendering at the viewpoints where the visible area spans multiple blocks. The Gaussians utilized here come from direct concatenation of fine-tuned Gaussians of corresponding blocks. As depicted in Fig. S2, rendering from a specific viewpoint may involve four or more blocks. Despite that, the rendered images exhibit no discernible discontinuities, showcasing smooth boundary transitions facilitated by our coarse global Gaussian prior, as discussed in Sec. 3.2 of our main paper.
@@
E Scene Manipulation
Figure S3: Illustration of city scene manipulation driven by explicit representation of CityGS. In Part (a), the dome of the original building in the first row is repainted to the desired color shown in the second row. In Part (b), the building of the first row is removed and replaced with the one shown in the second row. In Part (c), the cars parked at locations shown in the left image are moved to the positions shown in the right image, so as to simulate the required traffic conditions. NeRF-based methods struggle to realize such manipulation.
For the implicit representation of NeRF-based methods, it is hard to explain the correspondence between network parameters and scene structure. However, since we can reconstruct the explicit city representation with relatively high geometric precision in CityGS, the geometric and appearance distribution can be manipulated as desired. The demos are shown in Fig. S3. The appearance of a specified part of a building can be transformed to the desired style. It is also possible to delete a building and replace it with another one. By placing cars or pedestrians, the pre-defined traffic conditions can be simulated in the city. These demos indicate potential real-time and interactive application of CityGS.